A water sample was alkaline to both phenolphthalein and methyl orange end point. 50 ml water sample required 20 ml N/50 H2SO4 for P end point and another 5 ml for complete neutralisation Calculate type and amount of alkalinity present.


1
Expert's answer
2021-03-02T01:05:28-0500

C(H2SO44) = 0.02 / 2 = 0.01 M

CO32- + H+ = HCO3 - (20 ml)

HCO3- + H+ = H2O + CO2 (5 ml)

There is a mistake in a question: the volume untill P point is for the first equotion. It equals to half of amount of CO32-. The second volume (untill M point) is due to another half of CO32- plus all amount of HCO3. So the second volume can't be bigger than the first one.